A chemist, has a 30% hydrochloric acid solution and a 55% hydrochloric acid solution. How many liters of each should she mix to get 320.313 liters of a hydrochloric acid solution with a 39% acid concentration?

Answer by stanbon(75887)

A chemist, has a 30% hydrochloric acid solution and a 55% hydrochloric acid solution. How many liters of each should she mix to get 320.313 liters of a hydrochloric acid solution with a 39% acid concentration?
-----
Equation:
acid + acid = acid
0.30x + 0.55(320.313-x) = 0.39*320.313
------
30x + 55*320.313 - 55x = 39*320.313
-----------
-25x = -16*320.313
------
x = 205 liters (amt of 30% solution needed)
320.313-x = 115.313 liters (amt of 55% solution needed)
------------
Cheers,
Stan H.
-------------